For the equation 3x2+px+3 = 0 , p>0, if one of the root is square of the other root then value of p is ?

Grade:12th Pass

1 Answers

ruchi yadav
10 years ago
let roots of the quadratic are a and a2

a2 + a = - p/ 3

a3 = 1 ; so a = 1

1 + 1 = -p/3

p = -6


Thank You
Ruchi
Askiitians faculty